Please help me in both questions, thank you very much :)QUESTION 3 Ben is a musician who has just finished his degree in music technology. He is currently deciding whether or not to accept a job in Brisbane as a music teacher at a local school, or focus on being a touring musician. He ill earn a wage of $35,000/year as a teacher in Brisbane, and $25,000/year as a touring musician based in Melbourne. Rent is $10,000lyear in Brisbane, but only $5,000/year in Melbourne. Assume touring accommodation costs are $0 as the tour provides free accommodation for musicians. However, if Ben decides to stay in Brisbane, he will spend $2,000/year on travelling to visit his friend in Melbourne Select the item from the list provided to make the following statements true: By taking the teacher job in Brisbane, 1. Accounting profit 2. -$2000/year Ben will earn an compared to touring of S3000 3. Total cost 4. Economic profit 5. Economic loss 6. A normal profit 7. $12000/year 8. $33000/year 9. Explicit cost Bens explicit costs for choosing to take the teaching job in Brisbane are For Ben to be indifferent, from an economists point of view between 10. $45000/year teaching and touring, he would need to 11. $1900 make option. when choosing either 12. Implicit cost

QUESTION 5 Australians are one of the highest per capita consumers of coffee in the world. Their coffee consumption has been increasing over the years. As a result, a large number of cafes have continued to open across Brisbane to cater for this increasing demand. Answer the following questions: a. Imagine the market is currently at long run equilibrium. Research has recently been published indicating coffee consumption may lead to increased incidents of cancer. This negative publicity about coffee would be expected to lead to a decline in demand for coffee. As a result, one could expect firms to X for Exit, E for Enter or R for Remain In. the market. Type b. If Marcos café was previously selling 300 coffees per day at a price of $5 per coffee, but is now only selling 230 coffees per day at a price of S4 per coffee. By how much has Marcos economic profit per day declined if his average total costs (which include opportunity cost) were previously $2 per coffee, and are now $3 per coffee? S per day. Answer to the nearest whole number (with no decimal places) c. If Bryans cafe has an average total cost (which includes opportunity cost) of $4.10 per coffee and an average variable cost of $3.90 per coffee, should he keep selling coffees in the short run if the price is $4 per coffee? for No, or M for need More information. Type Y for Yes, N
Answer #1

Answer:

Question 4)

A) Economic Profit of $3000.

Since accounting profit= Total Salary from teaching job- Explicit cost

Normal profit = Total salary from From teaching -Total actual expenditure for living in brisbane-Implicit cost

Implicit cost=Total salary from From touring -Total actual expenditure for living in melbourne=25000-5000=20000

Normal profit =35000-(10000+2000)-20000=$3000

Since normal Profit is greater than Zero so it is economic profit.

B) $12000/year

Explicit cost for teaching job= rental Expenses+ travelling Expenses=10000+2000=$12000/year

c) Normal Profit

Normal Profit occur when Total salary = Total explicit cost+ Implicit cost and

which means Normal Profit =0 so Ben will be indifferent while choosing job.